OpenStudy (anonymous):

Change –5x + 2y = 12 into slope-intercept form.

OpenStudy (anonymous):

separate the y and x y = a(x) +- b does that help?

OpenStudy (anonymous):

sorry no

OpenStudy (anonymous):

-5x + 2y = 12 2y = 12 + 5x y = 12/2 + 5x/2 y = 6 + (5/2)x

OpenStudy (anonymous):

better now? hehe

OpenStudy (anonymous):

yes ty it makes more since

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Select the correct graph of –x – 2y = 6.